The base of an isosceles triangle is 30cm less than one of its sides. If the perimeter is 210cm find the dimension of the triangle.
Answer by lwsshak3(11628)

The base of an isosceles triangle is 30cm less than one of its sides. If the perimeter is 210cm find the dimension of the triangle.
***
let x=length of each of two sides
x-30=length of base
2x+x-30=210
3x=240
x=80
x-30=50
length of each of two sides=80 cm
length of base=50 cm